# Nonlinear Causality in Brain Networks: With Application to Motor Imagery vs Execution

### v1

@@ 2.4 On the role of TAR models for exploring causal connectivity in a network of multiple nodes
    Introducing TAR models into causality analysis for networks extends their applicability to interconnected systems. This approach offers researchers new perspectives to explore complex interactions, highlighting threshold nonlinear causality in dynamic environments. Unlike conventional Granger causality tests, which rely on linear assumptions, the use of TAR models allows for the detection of nonlinear causal relationships between variables. This is particularly important as many real-world relationships exhibit nonlinearities, and detecting such relationships enhances the scope of causality analysis. The inclusion of a threshold nonlinearity test helps identify points where causal connections between nodes change qualitatively, enhancing interpretability in dynamic networks like EEG. Identifying regime-switching dynamics adds a layer of interpretability about the nature of causal impact in a dynamic network such as the EEG setting and can offer insights into the conditions under which, for example, the node $X$ influences the node $Y$.
    TAR models accommodate various regime-specific dynamics, offering a more refined representation of causal interactions under different conditions, where linear models may fall short. This flexibility allows experts to tailor the model to the data, improving representation of the underlying causal mechanisms. However, TAR models can be more complex, potentially leading to over-fitting or over-parameterization, especially with limited data. Estimating TAR models, particularly identifying thresholds, can also be computationally intensive, posing challenges for large networks or datasets.
    3 Application to Real Data
@@ 3.1 Dataset and the experimental procedure
    In this section, we exemplify the utility of the proposed approach for identifying connectivities and causal interactions in a complex time series network by implementing it on a real-world EEG dataset. The dataset used in the analysis originated from an experiment that entailed a series of motor/imagery tasks, providing rich information on the temporal dynamics of brain activity. The whole experiment consisted of EEG recordings from 109 volunteers using a 64-channel setup and publicly available at PhysioNet.org (Schalk et al., 2004). The experimental protocol comprised 14 experimental runs for each subject, including two baseline runs (one with eyes open and one with eyes closed) and twelve task-oriented runs. These tasks required participants to either physically perform or imagine movements in response to visual cues on a screen, such as opening and closing fists or feet, depending on the position of a target. The tasks are divided into four categories: physical movement of opening and closing of the left-hand or right-hand fist, imaginary movement of opening and closing the left-hand or right-hand fist, physical movement of opening and closing of both fists’ or both feet’, and imaginary movement of opening and closing of both fists’ or both feet’. EEG data were sampled at 160HZ, along with an annotation channel indicating rest or the onset of motion (real or imagined) for specific movements. The EEG montage followed the international 10-10 system, omitting certain electrodes (i.e., Nz, F9, F10, FT9, FT10, A1, A2, TP9, TP10, P9, and P10).
    All available data underwent the pre-processing phase to filter out artifacts (i.e., non-brain related activities) only, such as eye blinks and muscle-related movements, thereby avoiding interference that would damage any temporal structure of the recordings. The pre-processing procedure consists of the following steps: Initially, the recordings are subjected to a high-pass filter at 0.5Hz to eliminate low-frequency noise. This is followed by the removal of AC power-related line noises at 60Hz. Next, Independent Component Analysis (ICA), facilitated by MNE-python routines (Gramfort et al., 2013; Larson et al., 2024), is implemented to isolate and exclude the source of artifact components. Finally, artifacts associated with eye blinking and muscle movements are identified and removed, ensuring the recordings are free from these specific types of interference.
    Despite all subjects’ recordings undergoing the specified standard pre-processing procedure, some structural problems, such as sawtooth trace, excessive blinking, broken channels, unbalanced epoching, etc., could not be resolved and were identified in some subjects’ records. Therefore, we limited the scope of our analysis to 35 Right-handed subjects (17 Females and 18 Males) whose recordings were found to be qualified. Further, to more appropriately demonstrate how the proposed method can be used to dissect causal connections between pairwise nodes within any network of time series, we focused on the associations among four nodes (i.e., channels), with two (i.e., AF3, AF4) positioned over the pre-frontal cortex and two (i.e., C3, C4) placed over the motor cortex illustrated in Fig. 1. These channels were chosen because they correspond to regions of the brain (i.e., prefrontal and motor cortices) that are likely to be active during voluntary movement and imagery.
    Figure 1: Channels of interest colored on a sample of 64-channel EEG placement.
    The connectivities between these channels were examined in 35 subjects’ data pertaining to both motor movement (MM) and motor imagery (MI) tasks. The subjects performed a series of opening and closing (OC) gestures, including left-hand fist (LH), right-hand fist (RH), both fists’ (BH), and both feet’ (BF), for both the motor movement (MM) and motor imagery (MI) tasks. The aim was to compare the brain activity associated with these tasks and gestures by identifying the patterns of connectivity and causal interactions amongst the selected channels. The connectivities explored can be outlined in Table 1.
    Table 1: Outline for illustrating causal connections during different brain activities.
    Motor Movement Motor Imagery
    LH [Uncaptioned image] [Uncaptioned image]
    RH [Uncaptioned image] [Uncaptioned image]
@@
    *The black arrows indicate the pairwise causal interactions explored.
    3.2 Implementation of the proposed approach
    Each subject repeated the relevant task (i.e., task1: OC of LH/RH, task2: imagery OC of LH/RH, task3: OC of BH/BF, task4: imagery OC of BH/BF) 3 times (i.e., runs). Within these three task-oriented runs, they performed each motor movement/imagery motion (i.e., LH, RH, BH, BF) approximately 23 times in total.
    Each epoch consists of about 4 seconds, that is, 640 observations. Initially, the relevant epochs were extracted from the dataset with the help of annotation information. For each epoch, 8 different causality interactions were tested pairwisely. For each task, 23 test results were combined using the method suggested by Becker (1994), and a decision was made on the presence of Threshold Non-linearity in the relevant task for the subject.
    The proposed method for inferring connectivity for just one pair of channels has been represented in Chapter 2 by the procedure named TAR4C. In the application to the whole dataset, we first focused on 2 index metrics in order to provide connectivity schemes covering the analysis results of all subjects considered.
@@ 3.2.1 Threshold Connectivity Index
    The Threshold Connectivity Index (TCI) quantifies, considering all the subjects, the presence of connectivity between two EEG channels based on the detection of threshold nonlinearity in their interaction. This index reflects the overall connectivity between the channels under specific conditions, such as during motor imagery or motor movement tasks.
@@
    where $n^{(d)}$ is the number of subjects out of the total $N$ who reject the linearity test at delay $d$. This indicates the number of subjects exhibiting threshold nonlinearity at the specified delay. $\mathbb{I}\left(n^{(d)}\geq 0.7\times N\right)$ is an indicator function. It equals 1 when $n^{(d)}\geq 0.7\times N$, meaning that the number of subjects rejecting linearity at delay $d$ equals or exceeds 70% of the total subjects. If this condition is not met, the indicator function equals 0. This ensures that only delays where a significant proportion of subjects (at least 70%) reject linearity are included in the TCI calculation. The resulting TCI value reflects the proportion of subjects showing significant threshold connectivity between channels $X$ and $Y$ and is expressed as a percentage. In practical terms, a higher TCI value suggests stronger threshold-based connectivity between the two channels, which is visualized in network diagrams with thicker lines representing stronger connections. The analysis procedure for calculating TCI involves the following steps:
@@
    Motor Movement / Motor Imagery
@@ 3.2.2 Threshold Granger Causality Index
    The Threshold Granger Causality Index (TGCI) extends the TCI by incorporating an additional step that tests the predictive power of the threshold variable. After establishing the presence of threshold nonlinearity, the TGCI assesses whether the past values of the threshold variable significantly predict the future values of the dependent variable using the Wald test. This index, therefore, provides a measure of directional causal influence between EEG channels.
    The TGCI is calculated as follows:
@@
    where $\text{Wald}^{(d)}$ is the number of subjects out of the total $N$ who pass the Wald test in favor of the predictive power of the threshold variable $X$ at delay $d$. This index quantifies the strength of the causal influence from $X$ to $Y$.
    The TGCI reflects the strength and directionality of the causal relationship between the channels and is used to visualize the directional flow of information in network diagrams. A higher TGCI value indicates a stronger causal influence from the threshold variable to the dependent variable.
    The analysis procedure for calculating TGCI builds on the steps used for TCI and includes the following additional steps:
@@
    Motor Movement / Motor Imagery
@@ 3.3 Connectivity at Motor Imagery
    MI left-hand fist gesture (MI-LH): For the MI-LH gesture, the TCI results (Table 4) indicate moderate connectivity between the channels. The connectivity from AF3 to C3 (TCI = 7.14, red arrow) and from AF4 to C4 (TCI = 7.86, red arrow) is slightly higher
    TCI at MI TGCI at MI
    LH
@@
    RH
    [Uncaptioned image] [Uncaptioned image]
    BHs
    [Uncaptioned image] [Uncaptioned image]
    BF
@@
    than the reverse connections (C3 to AF3, TCI = 6.67, blue arrow; C4 to AF4, TCI = 7.62, blue arrow). For AF4, the TCI values show 6.19 (red arrow) from AF4 to C3 and 6.43 (blue arrow) from C3 to AF4, indicating nearly equal connectivity in both directions. These results suggest somewhat balanced connectivity between these channels, with a slight prominence of the motor cortex over the prefrontal cortex in the C4 to AF3 (TCI = 14.29, blue arrow).
    The TGCI values for the MI-LH gesture (Table 4) show a more complex pattern. While the TCI indicated moderate connectivity, the TGCI values reveal that the causal influence from C4 to AF3 (TGCI = 12.62, blue arrow) is stronger than its reverse direction from AF3 to C4 (TGCI = 5.95, red arrow), and the causal influence from AF4 to C4 is notably high (TGCI = 25.24, red arrow) than its opposing direction from C4 to AF4 (TGCI = 13.1, red arrow). These indicate a significant role of AF4 and C4 in driving motor activity during left-hand fist motor imagery.
    MI right-hand fist gesture (MI-RH): The MI-RH gesture results exhibit stronger overall connectivity, as seen in the TCI values(Table 4). The connection from AF3 to C3 has a TCI of 13.10 (red arrow), while the reverse direction, C3 to AF3, has a TCI of 5.95 (blue arrow), indicating a stronger connectivity from AF3 over C3. The connection from AF3 to C4 has a TCI of 14.29 (red arrow), with the reverse, C4 to AF3, showing 6.67 (blue arrow), again highlighting stronger prefrontal influence. The connections involving AF4 also show similar patterns, with TCI values of 13.33 (red arrow) from AF4 to C3 and 6.19 (blue arrow) in the reverse direction, and 13.57 (red arrow) from AF4 to C4 with 12.62 (blue arrow) in the reverse.
    The TGCI values for MI-RH (Table 4) also reflect strong directional connectivity, particularly from the channels over the prefrontal to the motor cortices. AF3 to C4 and AF4 to C4 both have TGCI values of 18.33 (red arrows), indicating significant causal influence from the channels over the prefrontal cortex to the motor regions. The reverse connections, C4 to AF3 and C4 to AF4, both have lower TGCI values of 6.19 (blue arrows), suggesting that while there is some feedback, it is much weaker compared to the prefrontal influence.
    MI both hands’ fists gesture (MI-BHs) For the MI-BHs gesture, the TCI results (Table 4) show balanced and strong connectivity between the channels over the prefrontal and the motor cortices. The connection from AF3 to C3 has a TCI of 7.38 (red arrow), with the reverse direction showing 6.90 (blue arrow). The connection from AF3 to C4 has a TCI of 14.05 (red arrow), while C4 to AF3 is 13.57 (blue arrow), indicating nearly equal connectivity in both directions. The connectivity involving AF4 is similarly strong and balanced, with TCI values of 14.05 (red arrow) from AF4 to C3 and 12.62 (blue arrow) in the reverse direction, and 13.81 (red arrow) from AF4 to C4 with 7.86 (blue arrow) in the reverse.
    The TGCI results for MI-BH reveal significant contralateral causal influence, particularly between C3 and AF4 (TGCI = 17.86, blue arrow) and AF3 to C4 (TGCI = 12.14, red arrow). The connection from C3 to AF4 and having no flow in the opposite direction of itself indicate a strong causal influence from the channel over the left motor cortex to the right prefrontal cortex.
    MI both feet gesture (MI-BF) The MI-BF gesture (Table 4) shows the highest TCI values, particularly for the connections between AF3 and C4. The connection from AF3 to C4 has a TCI of 21.19 (red arrow), with the reverse, C4 to AF3, showing 20.95 (blue arrow), indicating exceptionally strong and nearly equal connectivity in both directions. The connections from AF4 to C3 and C4 also show high TCI values, with 13.10 (red arrow) from AF4 to C3 and 13.81 (blue arrow) in the reverse direction, and 14.76 (red arrow) from AF4 to C4 with 14.52 (blue arrow) in the reverse, indicating strong and balanced connectivity.
@@
    The TGCI results further emphasize this strong connectivity, with particularly high values for the feedback from C3 to AF4 (TGCI = 24.05, blue arrow) and from C4 to AF4 (TGCI = 18.10, blue arrow). These findings indicate that during both feet’ motor imagery, there is a continuous exchange of information between the channels over motor and prefrontal cortices, with the motor regions having a significant causal influence over the prefrontal areas. This strong bidirectional connectivity is crucial for the precise coordination required for bilateral foot movements.
    Comparative Analysis Between gestures:
    When comparing the MI-LH and MI-RH gestures, the RH gesture exhibits stronger overall connectivity. The TCI values for the RH gesture are higher across nearly all channel pairs compared to the LH gesture. For instance, the TCI from AF3 to C3 is 13.10 (red arrow) in the RH task compared to 7.14 (red arrow) in the LH task. Similarly, the connectivity from AF3 to C4 is stronger in the RH task (TCI = 14.29, red arrow) than in the LH task (TCI = 6.90, red arrow). This suggests that the RH motor imagery, which typically involves the dominant hand in most subjects, engages more robust neural connectivity, particularly from the channels over the prefrontal cortex to the motor cortices.
    The TGCI results further emphasize this difference. The MI-RH shows strong directional causal influence from AF3 and AF4 to C4, with TGCI values of 18.33 (red arrows) for both connections, indicating that the channels over the prefrontal cortex hold significant causal control over the motor cortex during this task. In contrast, the MI-LH shows weaker causal connectivity, with the most significant TGCI being 25.24 (red arrow) from AF4 to C4, but less pronounced overall compared to the RH task. This difference likely reflects the increased cognitive demand and neural resources required for controlling the dominant hand during motor imagery.
    The MI-BF gesture shows the highest TCI values, particularly in the interactions between AF3 and C4, and between C4 and AF3. This suggests that the motor imagery of both feet’ gesture involves stronger connectivity compared to other gestures, likely due to the increased complexity of coordinating bilateral foot movements.
    Across all gestures, a few general trends can be observed. First, bilateral tasks (BHs and BF) generally involve stronger and more complex channel connectivity compared to unilateral tasks (LH and RH). This is reflected in both the TCI and TGCI values, where bilateral tasks show higher indices, indicating more extensive communication and coordination between channels. This increased connectivity likely reflects the higher cognitive demands associated with coordinating movements across both sides of the body.
    Second, channels over the prefrontal cortex appear to play a more noticeable role in controlling channels over motor regions during right-hand fist tasks, as evidenced by the higher TCI and TGCI values for connections from AF3 and AF4 to C4 in the right-hand fist gesture. This prominence is less pronounced in the left-hand fist gesture, where connectivity and causal influences are generally weaker.
    Finally, the bidirectional connectivity observed in TGCI results, particularly during both fists’ and both feet’ gestures, highlights the significance of feedback between the channels over motor and over prefrontal regions during complex motor imagery tasks.
    3.4 Connectivity at Motor Movement
    MM left-hand fist gesture (MM-LH): For the MM-LH, the TCI results (Table 5) indicate pronounced connectivity from the channels over motor cortex back to the channels over prefrontal cortex, as evidenced by the high TCI value of 18.81 from C3 to AF3 (blue arrow). This is significantly higher than the connectivity from AF3 to C3 (TCI = 7.14, red arrow), suggesting that during actual left-hand fist movement, channel C3 maintains stronger connectivity on channel AF3 rather than the reverse. Similar patterns are observed between AF3 and C4, with a TCI of 13.81 (red arrow) from AF3 to C4 and a slightly lower TCI of 12.14 (blue arrow) from C4 to AF3. The connectivity between AF4 and C3/C4 is more balanced, with nearly equal TCI values in both directions.
    The TGCI results for MM-LH further highlight this motor-to-prefrontal influence, with a strikingly high TGCI of 25.00 (blue arrow) from C3 to AF3, indicating a strong causal influence. The causal influence from the prefrontal cortex to the motor regions is relatively weaker, as shown by the TGCI values of 5.95 (red arrow) from AF4 to C3. This suggests that during left-hand fist movement, the channels over the motor cortex are highly connected to and causally influential over the channels AF3 and AF4, likely reflecting the real-time processing required for executing the movement.
    TCI at MM TGCI at MM
    Left Fist

### v2

@@ 2.4 On the role of TAR models for exploring causal connectivity in a network of multiple nodes
    Unlike conventional Granger causality tests, which rely on linear assumptions, the use of TAR models allows for the detection of nonlinear causal relationships. This is particularly important as many real-world relationships exhibit nonlinearities, and detecting such relationships enhances the scope of causality analysis. The inclusion of a threshold nonlinearity test helps identify points where causal connections between nodes change qualitatively, enhancing interpretability in dynamic networks like EEG. Identifying regime-switching dynamics adds a layer of interpretability about the nature of causal impact in a dynamic network such as the EEG setting and can offer insights into the conditions under which, for example, the node $X$ influences the node $Y$. TAR models accommodate regime-specific dynamics, offering a refined representation of causal interactions where linear models may fail. This flexibility allows experts to better model the underlying causal mechanisms. However, TAR models can be more complex, leading to overfitting or over-parameterization, especially with limited data. Estimating TAR models and identifying thresholds can also be computationally intensive, posing challenges for large networks or datasets.
    3 Analysis of Motor Imagery vs Execution EEG Data
@@ 3.1 Dataset and the experimental procedure
    This paper examines differences in brain network dynamics between motor imagery (MI) and execution (ME) tasks. The experiment collected EEG recordings from 109 volunteers with a 64-channel setup, available publicly at PhysioNet.org (Schalk et al., 2004). Participants were asked to either execute or imagine motor tasks based on visual cues. Subjects performed gestures such as opening and closing (OC) their both fists/feet (BH/BF). EEG data were sampled at 160 Hz, with annotations for rest and activity (real or imagined). Data preprocessing removed artifacts like eye blinks and muscle activity to maintain the integrity of temporal structures. The steps included applying a high-pass filter at 0.5Hz, removing 60Hz line noise, and performing Independent Component Analysis (ICA) to isolate and eliminate artifact sources. Eye and muscle-related artifacts were identified and excluded. Despite the preprocessing, some issues (e.g., sawtooth traces, excessive blinking) remained in some recordings. We narrowed the analysis to 35 right-handed subjects (17 females, 18 males) with clean recordings. To focus on causal connections, we selected four channels (AF3, AF4, C3, C4) positioned over the prefrontal and motor cortices (Figure 1), which are likely to be active regions during voluntary movement and imagery. The connectivities explored can be outlined in Table 1.
    [Uncaptioned image] Figure 1: Channels of interest.
    Table 1: Illustration of explored connections.
    Motor Execution Motor Imagery
@@
    3.2 Analysis using the TAR4C model
    Each subject repeated the relevant task (i.e., task1: OC of BH/BF, task2: imagery OC of BH/BF) 3 times (i.e., runs). Within these three task-oriented runs, they performed each motor execution/imagery motion (i.e., BH, BF) approximately 23 times in total. Each epoch contains about 4 seconds (i.e., 640 observations).
    For each task, 23 test results were combined following Becker (1994) to determine the presence of threshold non-linearity in the task for each subject. The method for inferring connectivity for a channel pair, called TAR4C, is described in Chapter 2. In the application step, we focused on two index metrics to summarize connectivity results for all subjects.
@@ 3.2.1 Threshold Connectivity Index
    The Threshold Connectivity Index (TCI) quantifies, considering all the subjects, condition-specific connectivity between two EEG channels based on the detection of threshold nonlinearity in their interaction.
@@
    where $n^{(d)}$ is the number of subjects out of the total $N$ who reject the linearity test at delay $d$. This indicates the number of subjects exhibiting threshold nonlinearity at the specified delay. $\mathbb{I}\left(n^{(d)}\geq 0.7\times N\right)$ is an indicator function; it equals 1 when $n^{(d)}\geq 0.7\times N$, meaning that the number of subjects rejecting linearity at delay $d$ equals or exceeds 70% of the total subjects.
    In practical terms, a higher TCI value suggests stronger threshold-based connectivity between the two channels, which is visualized in network diagrams with thicker lines representing stronger connections. The analysis procedure for calculating TCI involves the following steps:
@@
    Motor Execution / Motor Imagery
@@ 3.2.2 Threshold Granger Causality Index
    The Threshold Granger Causality Index (TGCI) extends the TCI by incorporating an additional step that tests the predictive power of the threshold variable. After establishing the presence of threshold nonlinearity, the TGCI assesses whether the past values of the threshold variable significantly predict the future values of the dependent variable using the Wald test. This index, therefore, provides a measure of directional causal influence between EEG channels. The TGCI is calculated as follows:
@@
    where $\text{Wald}^{(d)}$ is the number of subjects out of the total $N$ who pass the Wald test in favor of the predictive power of the threshold variable $X$ at delay $d$. This index quantifies the strength of the causal influence from $X$to $Y$. The TGCI reflects the strength and directionality of the causal relationship between the channels and is used to visualize the directional flow of information in network diagrams. A higher TGCI value indicates a stronger causal influence from the threshold variable to the dependent variable. The analysis procedure for calculating TGCI builds on the steps used for TCI and includes the following additional steps:
@@
    Motor Execution / Motor Imagery
@@ 3.3 Connectivity at Motor Imagery
    MI both hands’ fists gesture (MI-BHs). For the MI-BHs gesture, the TCI results (Table 4) show balanced and strong connectivity between the channels over the prefrontal and the motor cortices. The TCI value for AF3 $\to$ C3 is 7.38 (red arrow), while C3 $\to$ AF3 is 6.90 (blue arrow). Meanwhile, AF3 $\to$ C4 is 14.05 (red arrow); C4 $\to$ AF3 is 13.57 (blue arrow), indicating nearly equal connectivity in both directions. The connectivity involving AF4 is similarly strong and balanced, with TCI value from AF4 $\to$ C3 to be 14.05 (red arrow) and C3 $\to$ AF4 to be 12.62 (blue arrow); AF4 $\to$ C4 is 13.81, C4 $\to$ AF4 is 7.86 (blue arrow). The TGCI values for MI-BH reveal significant contralateral causal influence in particular channels: C3 $\to$ AF4 17.86 (blue arrow), AF3 $\to$ C4 12.14 (red arrow). The connection from C3 $\to$ AF4 and having no flow in the opposite direction of itself indicate a strong causal influence from the channel over the left motor cortex to the right prefrontal cortex.
    Both Fists (BHs) Both Feet (BF)
    TCI at MI
@@
    TGCI at MI
@@
    MI both feet gesture (MI-BF). The MI-BF gesture (Table 4) shows the highest TCI values, particularly for AF3 $\leftrightarrow$ C4: AF3 $\to$ C4 21.19 (red arrow), C4 $\to$ AF3, 20.95 (blue arrow), indicating exceptionally strong and nearly equal connectivity in both directions. The connection AF4 $\leftrightarrow$ C3 and C4 also show high TCI values: AF4 $\to$ C3 13.10 (red arrow) and C3 $\to$ AF4 13.81 (blue arrow); AF4 $\to$ C4 14.76 (red arrow), C4 $\to$ AF4 14.52 (blue arrow), indicating strong and balanced connectivity. The TGCI results further emphasize this strong connectivity, with particularly high values for the C3 $\to$ AF4 24.05 (blue arrow) and C4 $\to$ AF4 18.10 (blue arrow). These findings indicate that during the imagery of both feet task, communication between the channels over the motor and prefrontal cortices continues, with the motor regions having a significant causal influence over the prefrontal areas. This prominent bidirectional interconnectivity appears to be an important element for the appropriate coordination of bilateral feet movement.
    Comparative analysis between gestures. The MI-BF gesture demonstrates stronger connectivity than MI-BHs, particularly between AF3 and C4. TCI values in MI-BF are higher (21.19 and 20.95 vs. 14.05 and 13.57 in BHs), reflecting more intense communication during MI-BF, with an increased role for the motor cortex. MI-BF highlights motor-to-prefrontal feedback, especially from [C3, C4] $\to$ AF4, pointing to stronger motor regulation of prefrontal activity during MI-BF. MI-BF also reveals more pronounced cross-hemispheric feedback between channels, suggesting denser bilateral engagement. Both tasks display strong bidirectional connectivity, with AF3 and AF4 affecting motor channels and vice versa. MI-BF suggests stronger motor-to-prefrontal feedback and overall higher connectivity, implying a more prominent dependence on motor cortex regulation during MI-BF.
    3.4 Connectivity at Motor Execution
    ME both hand’s fists gesture (ME-BHs). For MI-BHs gesture, the TCI results (Table 5) show relatively proportional connectivity across the prefrontal and motor cortices. The TCI values from AF4 $\to$ C3 is 14.29 (red arrow), and from AF3 $\to$ C4, it is 13.33 (red arrow), suggesting noteworthy connectivity over C3 and C4. However, the TGCI results for MI-BHs gesture seem to oppose this finding, with significant causal influence observed from C4 $\to$ AF4 12.14 (blue arrow) and from C3 $\to$ AF3 11.90 (blue arrow). The contrast in terms of TCI and TGCI during MI-BHs task suggests there is a continuous exchange of information between the channels, which might be related to coordinating bilateral motor activity. The bidirectional causal connectivity indicates that both hemispheres are actively engaged, with the prefrontal and motor cortices contributing to the planning and execution of the movement.
    Both Fists (BHs) Both Feet (BF)
    TCI at ME